OpenStudy (anonymous):

For an amortized loan, you are given the loan, annual interest rate, and length of the loan. Find the monthly payment necessary to pay off the loan. Do not round until the final answer. Then, round to the nearest cent. amount: $9,000; rate: 9.6%; time: 3 years A $286.13 B $298.74 C $285.49 D $288.72

OpenStudy (anonymous):

I know that the formula has to be something like this PMT = PVr /(1- (1+r)^-n) But i m confused ..

OpenStudy (anonymous):

Monthly Payment: = 9,000 * ( .096/12) / { 1 - [1+ ( .096/12)^ -36 ] } = $288.72